OpenStudy (anonymous):

the press box at a baseball park is 33.0 ft above the ground. A reporter in the press box looks at an angle of 19.0 degrees below the horizontal to see second base. what is the horizontal distance form the press box to second base? This is what i did.... 90+19.0=109 180-109=71 (sin 90x)(sin 19/33) =.4867547088 I THINK I'AM WRONG. CAN SOMEONE HELP ME! PLEASE. THANKS

OpenStudy (anonymous):

my question is how to do this equation about press box

OpenStudy (anonymous):

The 33 ft is the vertical distance on the right angle triangle and you are trying to find the horizontal distance. So you can do: tan 19.0 = 33.0 / x x =33/tan19.0 which equals roughly 95.8 ft

OpenStudy (anonymous):

thank you)

OpenStudy (anonymous):